# Communication-Efficient Distributed Deep Learning via Federated Dynamic Averaging

### v2

@@ Abstract.
Driven by the ever-growing volume and decentralized nature of data, coupled with the need to harness this data and generate knowledge from it, has led to the extensive use of distributed deep learning (DDL) techniques for training. These techniques rely on local training that is performed at the distributed nodes based on locally collected data, followed by a periodic synchronization process that combines these models to create a global model. However, frequent synchronization of DL models, encompassing millions to many billions of parameters, creates a communication bottleneck, severely hindering scalability. Worse yet, DDL algorithms typically waste valuable bandwidth, and make themselves less practical in bandwidth-constrained federated settings, by relying on overly simplistic, periodic, and rigid synchronization schedules. These drawbacks also have a direct impact on the time required for the training process, necessitating excessive time for data communication. To address these shortcomings, we propose Federated Dynamic Averaging (FDA), a communication-efficient DDL strategy that dynamically triggers synchronization based on the value of the model variance. In essence, the costly synchronization step is triggered only if the local models, which are initialized from a common global model after each synchronization, have significantly diverged. This decision is facilitated by the communication of a small local state from each distributed node/worker. Through extensive experiments across a wide range of learning tasks we demonstrate that FDA reduces communication cost by orders of magnitude, compared to both traditional and cutting-edge communication-efficient algorithms. Additionally, we show that FDA maintains robust performance across diverse data heterogeneity settings.
Federated Learning, Distributed Learning
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 26th International Conference on Extending Database Technology; March 25–28, 2025; Barcelona, Spainisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learningccs: Computing methodologies Distributed artificial intelligenceccs: Computing methodologies Parallel computing methodologies
@@ 1. Introduction
The big data era has been marked by an unprecedented scale of training datasets (Zhang et al., 2021; Rauf et al., 2024). These datasets are not only growing in size, but are often physically distributed and cannot be easily centralized due to business considerations, privacy concerns, bandwidth limitations (especially in federated settings, such as drones collecting and collaboratively building a global model/view of an area), and data sovereignty laws (Kairouz et al., 2021; Wu et al., 2020; Cormode et al., 2024). Such constraints complicate the use of Deep Learning (DL) techniques in the aforementioned scenarios, and provide challenges in harnessing knowledge from distributed data.
Distributed Deep Learning (DDL) has emerged as an alternative paradigm to the traditional centralized approach (Zinkevich et al., 2010; Chilimbi et al., 2014), offering efficient handling of large-scale data across multiple worker-nodes, enhancing the speed of training DL models and paving the way for more scalable and resilient DL applications (Um et al., 2023; Ma et al., 2023; Lai et al., 2023; Zhou et al., 2022; Davitkova et al., 2024). Most DDL methods are iterative, where, in each iteration, some amount of local training is followed by synchronization of the local models with the global one. The predominant method, based on the bulk synchronous parallel (BSP) approach (Valiant, 1990), is to average the local model updates and then apply the average update to each local model (Zinkevich et al., 2010). Less synchronized variants have also been proposed, to ameliorate the effect of straggler workers (Miao et al., 2021; Fu et al., 2022) but compromise convergence speed and model quality.
A significant challenge inherent in the traditional techniques, especially in federated DL settings, where models are huge and worker interconnections are slow, is the communication bottleneck, restricting system scalability (Tang et al., 2023; Wang et al., 2016). This leads to a low computation-to-communication ratio (Shi et al., 2021; Fu et al., 2022). Addressing the communication overhead to expedite DDL algorithms has been a focal point of research for many years; speeding-up SGD is arguably among the most impactful and transformative problems in machine learning (Wang and Joshi, 2021).
@@
FDA utilizes a novel 2-action, conditional synchronization protocol, designed to avoid the need to decide or guess the proper values of local update steps, or to synchronize after each training step, but rather only performs the costly synchronization process when needed. Our FDA algorithm dynamically triggers synchronization based on the value of model variance across nodes. In a nutshell, the costly synchronization step is only triggered if the local models have diverged significantly, which implies that the global model may no longer be accurate.
As Figure 1 demonstrates, at the start, nodes enter the local training step with the same global model (Figure 1.(A)). Then, local training commences and each distributed node/worker computes its local state (Figure 1.(B)). This is followed by the transmission (Figure 1.(C)) of these small-size local states, an operation that is bandwidth- and time-efficient because of their small size. During transmission, the local states are aggregated and their average is made available to all nodes—an operation known as AllReduce. This operation does not require (or prohibit) the use of a central node. Based on the aggregated state, the nodes can estimate (Figure 1.(D)) whether the variance of the local models may have exceeded a threshold. If this is not the case, the costly synchronization step (Figure 1.(E)) is avoided and local training continues. What is important is how to properly pick these local states computed at, and then transmitted by, the local nodes. To address this problem, we propose two variants of our FDA algorithm. Our contributions can be summarized as follows:
  • We propose FDA, an algorithm that dynamically decides to synchronize local nodes when model variance across nodes exceeds a threshold. This strategy drastically reduces communication requirements, while preserving cohesive progress towards the shared training objective.
  • We propose two variants of FDA, which differ on the local state that is computed at, and then transmitted by, each distributed node/worker. These two variants are termed SketchFDA and LinearFDA.
  • We evaluate and compare FDA with other DDL algorithms through a comprehensive suite of experiments with diverse datasets, models, and tasks. Our experiments demonstrate that FDA outperforms traditional and contemporary FL algorithms by orders of magnitude in communication, while maintaining equivalent model performance. Furthermore, it effectively balances the competing demands of communication and computation, providing greatly improved trade-offs.
  • We show that FDA remains robust and efficient under various data distributions (Non-IID settings), maintaining comparable performance to the IID case. We note that in Federated Learning scenarios, the norm is that data are not independently and identically distributed across nodes.
@@
Figure 1. FDA. The local training step is followed by the computation of a local state by all nodes. Then, the (small in size) local states are aggregated. Based on the aggregated result, all nodes estimate if synchronization is required. In most cases, the expensive synchronization step of the models is avoided and local training continues
@@ 2. Related Work
Problem formulation. Consider distributed training of deep neural networks over multiple workers (Dean et al., 2012; Li et al., 2021). In this setting, each worker has access to its own set of training data $\mathcal{D}_{k}$, and the collective goal is to find a common model $\mathbf{w}\in\mathbb{R}^{d}$ by minimizing the overall training loss. This scenario can be effectively modeled as a distributed optimization problem, formulated as follows:
@@
Communication efficient Local-SGD. The work in (Li et al., 2021) decomposes each round into two phases. In the first phase, each worker runs Local-SGD with $\tau=I_{1}$, while the second phase run $I_{2}$ steps with $\tau=1$(Li et al., 2021) proposes to exponentially decay $I_{1}$ every $M$ rounds. In the heterogeneous setting, the work in (Qin et al., 2023), by analysing the convergence rate, proposes an increasing sequence of local update steps for strongly-convex local objectives, and suggests fixed local steps for convex, non-convex local objectives. The study in (Yu and Jin, 2019) dynamically increases batch sizes to reduce communication rounds, maintaining the same convergence rate as SSP-SGD. However, the large-batch approach leads to poor generalization (Hoffer et al., 2017), a challenge addressed by the post-local SGD method (Lin et al., 2020), which divides training into two phases: BSP-SGD followed by Local-SGD with a fixed number of steps. In the Lazily Aggregated Algorithm (LAG) (Chen et al., 2018), a different approach was taken, using only new gradients from some selected workers and reusing the outdated gradients from the rest, which essentially skips communication rounds.
@@
Federated Averaging (FedAvg) (McMahan et al., 2017) is another representative communication efficient Local-SGD algorithm, which is a pivotal method in Federated Learning (FL) (Kairouz et al., 2021). In the FL setting with edge computing systems, the work in (Wang et al., 2019) tries to find the optimal synchronization period $\tau$ subject to local computation and aggregation constraints. Recently (Mills et al., 2023), in the FL setting with the assumption of strongly-convex objectives, by analysing the balance between fast convergence and higher-round completion rate, a decaying local update step scheme emerged.
@@
Accelerating convergence. An indirect, yet highly effective, way to mitigate the communication burden in DDL, is to speed up convergence. Consequently, recent works have deployed accelerated versions of SGD to the distributed setting. Specifically, FedAdam (Reddi et al., 2021) extends Adam (Kingma and Ba, 2015) and FedAvgM (Hsu et al., 2019) extends SGD with momentum (SGD-M) (Sutskever et al., 2013a). Recently, Mime (Karimireddy et al., 2021) provides a framework to adapt arbitrary centralized optimization algorithms to the FL setting. However, these methods still suffer from the model divergence problem, particularly in heterogeneous settings. When solving (1), the disparity between each worker’s optimal solution $\mathbf{w}_{k}^{*}$ for their objective $F_{k}$, and the global optimum $\mathbf{w}^{*}$ for $F$, can potentially cause worker models to diverge (drift) towards their disparate minima (Karimireddy et al., 2020; Reddi et al., 2021; Wenig and Papenbrock, 2022). The result is slow and unstable convergence with significant communication overhead. To address this problem, the SCAFFOLD algorithm (Karimireddy et al., 2020) used control-variates (in the same spirit to SVRG), with significant speed-up. FedProx (Sahu et al., 2018) re-parameterized FedAvg (McMahan et al., 2017) by adding $L^{2}$ regularization in the workers’ objectives to be near the global model. Lastly, FedDyn (Acar et al., 2021) improved upon these ideas with a dynamic regularizer making sure that if local models converge to a consensus, this consensus point aligns with the stationary point of the global objective function.
@@
Compression. To reduce communication overhead in DDL, significant efforts have been directed towards minimizing message sizes. Key strategies include sparsification, where only crucial components of information are transmitted, as explored in (Aji and Heafield, 2017), and quantization techniques, which involve transmitting only quantized gradients, as detailed in (Shlezinger et al., 2021). Notably, these compression techniques can be combined with Local-SGD methods to enhance communication-efficiency further. A prime example is Qsparse-local-SGD (Basu et al., 2019), which integrates aggressive sparsification and quantization with Local-SGD, achieving substantial communication savings to reach specific target accuracies.
@@ 3. Federated Dynamic Averaging
Furthermore, let $\textsc{Optimize}(\mathbf{w},\mathcal{B})$ be the updated model vector (Goodfellow et al., 2016) computed by some optimization algorithm (e.g., SGD, Adam) using the model $\mathbf{w}$, and the batch $\mathcal{B}$ of training data. It incorporates the learning rate, loss function and relevant gradients. During step $t$, each worker $k$ first applies the update:
@@
In prior work, workers may (or not) synchronize by executing $\textsc{AllReduce}(\mathbf{w}_{t}^{(k)})$, thereby setting $\mathbf{w}_{t}^{(k)}=\overline{\mathbf{w}}_{t}$. If synchronization is not performed at step $t$, then each worker continues with the locally updated model.
@@
$\overline{\mathbf{w}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{w}_{t}^{(k)}$ Average model vector
@@
$\mathbf{\sigma}_{t}^{(k)}$ Lower-dim. proxy of $\mathbf{u}_{t}^{(k)}$
$\mathbf{S}_{t}^{(k)}=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\mathbf{\sigma}_{t}^{(k)}\right)$ State of worker $k$
@@
To preserve the RI, our FDA algorithm maintains (Lines 4-6 of Algorithm 1) at each node $k$ a local (low-dimensional) state-vector $\mathbf{S}_{t}^{(k)}$, which is computed based on $\mathbf{w}_{t}^{(k)}$. These state vectors are vital for the subsequent estimation of the model variance, and underpin the two variants of the FDA algorithm (provided in Sections 3.1 and 3.2, respectively). Our estimation techniques begin by performing AllReduce on the states $\mathbf{S}_{t}^{(k)}$, consolidating them into the average state $\overline{\mathbf{S}}_{t}$ (Line 7). Importantly, this communication step requires significantly less bandwidth and resources than transmitting the full models $\mathbf{w}^{(k)}_{t}$.
@@
Efficiently Monitoring the RI. Estimating model variance efficiently is at the heart of FDA. To this end, we first introduce the local model drift
$\mathbf{u}_{t}^{(k)}=\mathbf{w}_{t}^{(k)}-\overline{\mathbf{w}}_{t_{0}}$
where $\overline{\mathbf{w}}_{t_{0}}$ is the model vector after the most recent synchronization. Subsequently, the model variance can be written as:
@@
Proof. Adding an offset ($-\overline{\mathbf{w}}_{t_{0}}$) to each $\mathbf{w}_{t}^{(k)}$ does not alter the variance, therefore:
@@
10:          Update $S_{t}^{(k)}$
11:          $\overline{\mathbf{S}}_{t}\leftarrow\textsc{AllReduce}(S_{t}^{(k)})$
@@
Equation (4) indicates that the model variance estimation is reduced to estimating $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ efficiently. Computing $\left\|\overline{\mathbf{u}}_{t}\right\|_{2}^{2}$ exactly can be done by an AllReduce operation, but this would incur a communication cost equal to performing synchronization, since the dimension of model drifts is equal to that of the models.
We next present two techniques for communication-efficient variance estimation. To present them in a uniform manner, we define the local state $\mathbf{S}_{t}^{(k)}$, and the average state $\overline{\mathbf{S}}_{t}$ as
$\mathbf{S}_{t}^{(k)}=\left(\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\mathbf{\sigma}_{t}^{(k)}\right)\quad,\quad\overline{\mathbf{S}}_{t}=\left(\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{u}_{t}^{(k)}\right\|_{2}^{2}\,,\,\frac{1}{K}\sum_{k=1}^{K}\mathbf{\sigma}_{t}^{(k)}\right)$
where vectors $\mathbf{\sigma}_{t}^{(k)}$ are lower-dimensional representations of $\mathbf{u}_{t}^{(k)}$.
In the two subsequent sections, we formally define $\mathbf{\sigma}_{t}^{(k)}$, and a function $H$ that is applied on the average state vector $\overline{\mathbf{S}}_{t}$. Each specification of $\mathbf{\sigma}_{t}^{(k)}$ and $H$ corresponds to a different FDA variant.

### v4

@@ Abstract.
The ever-growing volume and decentralized nature of data, coupled with the need to harness it and extract knowledge, have led to the extensive use of distributed deep learning (DDL) techniques for training. These techniques rely on local training performed at distributed nodes using locally collected data, followed by a periodic synchronization process that combines these models to create a unified global model. However, the frequent synchronization of deep learning models, encompassing millions to many billions of parameters, creates a communication bottleneck, severely hindering scalability. Worse yet, DDL algorithms typically waste valuable bandwidth and render themselves less practical in bandwidth-constrained federated settings by relying on overly simplistic, periodic, and rigid synchronization schedules. These inefficiencies make the training process increasingly impractical as they demand excessive time for data communication. To address these shortcomings, we propose Federated Dynamic Averaging (FDA), a communication-efficient DDL strategy that dynamically triggers synchronization based on the value of the model variance. In essence, the costly synchronization step is triggered only if the local models—initialized from a common global model after each synchronization—have significantly diverged. This decision is facilitated by the transmission of a small local state from each distributed node. Through extensive experiments across a wide range of learning tasks we demonstrate that FDA reduces communication cost by orders of magnitude, compared to both traditional and cutting-edge communication-efficient algorithms. Additionally, we show that FDA maintains robust performance across diverse data heterogeneity settings.
copyright: rightsretaineddoi: isbn: 978-3-89318-098-1conference: 28th International Conference on Extending Database Technology (EDBT); 25th March-28th March, 2025; Barcelona, Spainjournalyear: 2025
@@ 1. Introduction
The big data era has been marked by an unprecedented scale of training datasets (Zhang et al., 2021; Rauf et al., 2024). These datasets are not only growing in size, but are often physically distributed and cannot be easily centralized due to business considerations, privacy concerns, bandwidth limitations—particularly in federated settings, such as drones collecting and collaboratively building a global model or view of an area—and data sovereignty laws (Kairouz et al., 2021; Wu et al., 2020; Cormode et al., 2024). Such constraints pose significant challenges to the application of Deep Learning (DL) techniques.
Distributed Deep Learning (DDL) has emerged as an alternative paradigm to the traditional centralized approach (Zinkevich et al., 2010; Chilimbi et al., 2014), offering efficient learning over large-scale data across multiple worker-nodes, enhancing the speed of training DL models and paving the way for more scalable and resilient DL applications (Um et al., 2023; Ma et al., 2023; Lai et al., 2023; Zhou et al., 2022; Davitkova et al., 2024). Most DDL methods are iterative, where each iteration involves a phase of local training followed by the synchronization of the local models with the global one.
The predominant method, based on the bulk synchronous parallel (BSP) approach (Valiant, 1990), averages the local model updates and applies the averaged update to each local model (Zinkevich et al., 2010). Less synchronized variants have also been proposed to ameliorate the effect of straggler workers, but they often compromise convergence speed and model quality (Miao et al., 2021; Fu et al., 2022).
A significant challenge inherent in the traditional techniques, especially in federated DL settings where models are huge and worker interconnections are slow, is the communication bottleneck, which restricts system scalability (Tang et al., 2023; Wang et al., 2016). Specifically, the communication bottleneck arises from the frequent exchange (synchronization) of model parameters—often in the range of billions—across distributed workers. The synchronization process entails substantial data volume transfer and generally dominates the overall training time, leading to a low computation-to-communication ratio (Shi et al., 2021; Fu et al., 2022). Addressing this challenge to expedite DDL algorithms has been a focal point of research for many years; speeding-up SGD is arguably among the most impactful and transformative problems in machine learning (Wang and Joshi, 2021).
@@
FDA utilizes a novel 2-action, conditional synchronization protocol, designed to avoid the need to decide or guess the proper values of local update steps, or to synchronize after each training step, but rather only performs the costly synchronization process when needed. Our FDA algorithm dynamically triggers synchronization based on the value of model variance across worker-nodes. In a nutshell, the costly synchronization step is only triggered if the local models have diverged significantly, which implies that the global model may no longer be accurate.
As Figure 1 demonstrates, at the start, workers enter the local training step with the same global model (Figure 1.A). Then, local training commences and each distributed worker-node computes its local state, which encapsulates helpful information for estimating the model variance (Figure 1.B). This is followed by the transmission (Figure 1.C) of these small-size local states, an operation that is bandwidth- and time-efficient because of their small size. During transmission, the local states are aggregated and their average is made available to all workers—an operation known as AllReduce. This operation does not require (or prohibit) the use of a central node. Based on the aggregated state, the workers can estimate (Figure 1.D) whether the variance of the local models may have exceeded a threshold. If this is not the case, the costly synchronization step (Figure 1.E) is avoided and local training continues. What is important is how to properly pick these local states computed at, and then transmitted by, the local workers. To address this problem, we propose two variants of our FDA algorithm. Our contributions can be summarized as follows:
  • We propose FDA, an algorithm that dynamically decides to synchronize local workers when model variance across workers exceeds a threshold. This strategy drastically reduces communication, while preserving cohesive progress towards the shared training objective.
  • We propose two variants of FDA, which differ in the amount of information preserved in the local states that are transmitted by each worker and aggregated for subsequent estimation of model variance. These two variants, termed SketchFDA and LinearFDA, offer a different balance between communication efficiency and approximation accuracy.
  • We evaluate and compare FDA with other DDL algorithms through a comprehensive suite of experiments with diverse datasets, models, and tasks. Our experiments demonstrate that FDA outperforms traditional and contemporary FL algorithms by 1-2 orders of magnitude in communication savings, while maintaining equivalent model performance. Furthermore, it effectively balances the competing demands of communication and computation, providing greatly improved trade-offs.
  • We demonstrate FDA’s robustness in various challenging Non-IID settings, common in real-world Federated Learning applications. While state-of-the-art methods typically require substantially more resources to converge under Non-IID conditions, FDA maintains consistent and comparable performance across both IID and Non-IID settings.
@@
Figure 1. FDA. The local training step is followed by the computation of a local state by all worker-nodes. Then, the (small in size) local states are aggregated. Based on the aggregated result, all workers estimate if synchronization is required. In most cases, the expensive synchronization step of the models is avoided and local training continues
@@ 2. Related Work
Problem formulation. Consider distributed training of deep neural networks over multiple workers (Dean et al., 2012; Li et al., 2021). In this setting, each worker represents a data owner (equivalently, a local model owner) and has access to its own set of training data $\mathcal{D}_{k}$. Workers can utilize any available hardware they possess (e.g., GPUs, CPUs) to perform learning steps. The collective goal is to find a common model $\mathbf{w}\in\mathbb{R}^{d}$ by minimizing the overall training loss. This scenario can be effectively modeled as a distributed optimization problem, formulated as follows:
@@
Solution direction. As noted in the seminal work (Kairouz et al., 2021), research in FL should focus primarily on synchronous solutions. This allows different lines of research (e.g., compression, privacy, etc.) to be developed independently and then combined seamlessly. Our work, along with most communication-efficient FL strategies, adheres to this synchronous paradigm. However, such approaches may be less effective in environments where each communication operation incurs significant overhead regardless of the size of the data being transmitted (e.g., high-latency). In these scenarios, asynchronous mechanisms become necessary, though they typically fall outside the primary focus of contemporary FL research. That said, FDA can be modified to work asynchronously (as explained in Section 3.3).
@@
Communication efficient Local-SGD. The work in (Li et al., 2021) decomposes each round into two phases. In the first phase, each worker runs Local-SGD with $\tau=I_{1}$, while the second phase runs $I_{2}$ steps with $\tau=1$(Li et al., 2021) proposes to exponentially decay $I_{1}$ every $M$ rounds. In the heterogeneous setting, the work in (Qin et al., 2023), by analysing the convergence rate, proposes an increasing sequence of local update steps for strongly-convex local objectives and fixed local update steps for other types of local objectives. The study in (Yu and Jin, 2019) dynamically increases batch sizes to reduce communication rounds, maintaining the same convergence rate as SSP-SGD. However, the large-batch approach leads to poor generalization (Hoffer et al., 2017), a challenge addressed by the post-local SGD method (Lin et al., 2020), which divides training into two phases: BSP-SGD followed by Local-SGD with a fixed number of steps. In the Lazily Aggregated Algorithm (LAG) (Chen et al., 2018), a different approach was taken, using only new gradients from some selected workers and reusing the outdated gradients from the rest, which essentially skips communication rounds.
@@
Federated Averaging (FedAvg) (McMahan et al., 2017) is another representative of communication efficient Local-SGD algorithms, which is a pivotal method in Federated Learning (FL) (Kairouz et al., 2021). In the FL setting with edge computing systems, the work in (Wang et al., 2019) tries to find the optimal synchronization period $\tau$ subject to local computation and aggregation constraints. Recently (Mills et al., 2023), in the FL setting with the assumption of strongly-convex objectives, by analysing the balance between fast convergence and higher-round completion rate, a decaying local update step scheme emerged.
Unlike previous approaches that rely on predetermined synchronization schedules (fixed, decaying, or otherwise), our work introduces a dynamic synchronization strategy. FDA adapts continuously during the training process, basing synchronization decisions on a real-time metric: the model variance across workers.
@@
Accelerating convergence. An indirect, yet highly effective way to mitigate the communication burden in DDL, is to speed up convergence. Consequently, recent works have built upon communication efficient Local-SGD methods by deploying accelerated versions of SGD to the distributed setting. Specifically, FedAdam (Reddi et al., 2021) extends Adam (Kingma and Ba, 2015) and FedAvgM (Hsu et al., 2019) extends SGD with momentum (SGD-M) (Sutskever et al., 2013a). Recently, Mime (Karimireddy et al., 2021) provides a framework to adapt arbitrary centralized optimization algorithms to the FL setting. However, these methods still suffer from the model divergence problem, particularly in heterogeneous settings. When solving (1), the disparity between each worker’s optimal solution $\mathbf{w}_{k}^{*}$ for their objective $F_{k}$, and the global optimum $\mathbf{w}^{*}$ for $F$, can potentially cause worker models to diverge (drift) towards their disparate minima (Karimireddy et al., 2020; Reddi et al., 2021; Wenig and Papenbrock, 2022). The result is slow and unstable convergence with significant communication overhead. To address this problem, the SCAFFOLD algorithm (Karimireddy et al., 2020) used control-variates (in the same spirit to SVRG), with significant speed-up. FedProx (Sahu et al., 2018) re-parameterized FedAvg (McMahan et al., 2017) by adding $L^{2}$ regularization in the workers’ objectives to be near the global model. Lastly, FedDyn (Acar et al., 2021) improved upon these ideas with a dynamic regularizer making sure that if local models converge to a consensus, this consensus point aligns with the stationary point of the global objective function.
While these approaches primarily focus on enhancing the optimization process and typically employ fixed synchronization intervals (e.g., every local epoch), our work addresses a complementary aspect: determining the optimal timing for synchronization. FDA’s dynamic synchronization strategy is orthogonal to these optimization techniques and can be integrated with them by simply adjusting the synchronization decision.
@@
Compression. To reduce communication overhead in DDL, significant efforts have been directed towards minimizing message sizes. Key strategies include sparsification, where only crucial components of information are transmitted, as explored in (Aji and Heafield, 2017), and quantization techniques, which involve transmitting only quantized gradients, as detailed in (Shlezinger et al., 2021). These techniques can be combined with Local-SGD methods to enhance communication-efficiency further. An example is Qsparse-local-SGD (Basu et al., 2019), which integrates aggressive sparsification and quantization with Local-SGD, achieving substantial communication savings. Crucially, FDA is fully compatible with any technique that reduces the cost of synchronization (e.g. model compression). Our approach simply adjusts the timing of the synchronization decision without altering the data being synchronized. This ensures that any compression technique effective in traditional methods (BSP, Local-SGD, etc.) will be equally effective when deployed with FDA. Therefore, the communication savings demonstrated in the relevant literature (Wang et al., 2023) can be safely expected to carry over to our approach as well.
@@ 3. Federated Dynamic Averaging
Furthermore, let $\textsc{Optimize}(\mathbf{w},\mathcal{B})$ be the updated model (Goodfellow et al., 2016) computed by some optimization algorithm (e.g., SGD, Adam) using the model $\mathbf{w}$, and the batch $\mathcal{B}$ of training data. It incorporates the learning rate, loss function and relevant gradients. During step $t$, each worker $k$ first applies the update:
@@
Workers synchronize by executing $\textsc{AllReduce}(\mathbf{w}_{t}^{(k)})$, thereby setting $\mathbf{w}_{t}^{(k)}:=\overline{\mathbf{w}}_{t}$. If synchronization is not performed at step $t$, each worker continues training with its locally updated model. A comprehensive list of the notation used throughout this section is provided in Table 1.
@@
$\mathcal{D}_{k}$ Training data of worker $k$
$\mathcal{B}^{(k)}_{t}$ A batch sampled from $\mathcal{D}_{k}$
@@
$\overline{\mathbf{w}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{w}_{t}^{(k)}$ Average model (global model)
@@
$\overline{\mathbf{w}}_{t_{-1}}$ Model after 2nd most recent sync.
@@
$\overline{\mathbf{u}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{u}_{t}^{(k)}$ Average model drift (global drift)
@@
$\mathbf{S}_{t}^{(k)}$ State of worker $k$
@@
$H(\cdot)$ Function for variance estimation
$\operatorname{sk}(\cdot):\mathbb{R}^{d}\to\mathbb{R}^{l\times m}$ AMS sketch operator (§3.1)
$\mathcal{M}_{2}(\cdot):\mathbb{R}^{l\times m}\to\mathbb{R}$ $L^{2}$ norm squared estimate (§3.1)
$\epsilon$ Error of sketch estimate (§3.1)
$(1-\delta)$ Confidence of approximation (§3.1)
$l=\mathcal{O}(\log 1/\delta)$Rows of sketch matrix (§3.1)
$m=\mathcal{O}(1/\epsilon^{2})$Columns of sketch matrix (§3.1)
$\xi=\frac{\overline{\mathbf{w}}_{t_{0}}-\overline{\mathbf{w}}_{t_{-1}}}{\left\|\overline{\mathbf{w}}_{t_{0}}-\overline{\mathbf{w}}_{t_{-1}}\right\|_{2}}$ Heuristic vec. for LinearFDA3.2)
@@
To preserve the RI, our FDA algorithm maintains (Lines 4-6 of Algorithm 1) at each worker $k$ a local (low-dimensional) state-vector $\mathbf{S}_{t}^{(k)}$, which is computed based on $\mathbf{w}_{t}^{(k)}$. These state vectors are vital for the subsequent estimation of the model variance, and underpin the two variants of the FDA algorithm (provided in Sections 3.1 and 3.2, respectively). Our estimation techniques begin by performing AllReduce on the states $\mathbf{S}_{t}^{(k)}$, consolidating them into the average state $\overline{\mathbf{S}}_{t}$ (Line 7). Importantly, this communication step requires significantly less bandwidth and resources than transmitting the full models $\mathbf{w}^{(k)}_{t}$.
@@
Efficiently Monitoring the RI. Estimating model variance efficiently is at the heart of FDA. To this end, we first introduce the local model drift, $\mathbf{u}_{t}^{(k)}$, and average drift, $\overline{\mathbf{u}}_{t}$, defined as follows:
$\mathbf{u}_{t}^{(k)}=\mathbf{w}_{t}^{(k)}-\overline{\mathbf{w}}_{t_{0}}\quad,\quad\overline{\mathbf{u}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{u}_{t}^{(k)}$
Here, $\overline{\mathbf{w}}_{t_{0}}$ denotes the model vector after the most recent synchronization. Subsequently, the model variance can be written as:
@@
Proof.
Adding an offset ($-\overline{\mathbf{w}}_{t_{0}}$) to each $\mathbf{w}_{t}^{(k)}$ does not alter the variance, therefore:
@@
10:          Update $\mathbf{S}_{t}^{(k)}$
11:          $\overline{\mathbf{S}}_{t}\leftarrow\textsc{AllReduce}(\mathbf{S}_{t}^{(k)})$
@@
Conceptually, following Eq (4), to precisely monitor the variance, we need to calculate two quantities: (1) $\frac{1}{K}\sum_{k=1}^{K}\|\mathbf{u}^{(k)}_{t}\|_{2}^{2}$, and (2) $\|\overline{\mathbf{u}}_{t}\|_{2}^{2}$. The first quantity requires an AllReduce operation on the squared norm of the worker drifts, which involves minimal overhead since these values are scalar. In contrast, the second quantity necessitates an AllReduce operation on the worker drifts themselves, which are of model dimension, thus incurring a high communication cost. In fact, this operation is equivalent to synchronization, which is exactly what we aim to avoid in the first place. Thus, it becomes evident that communication-efficient model variance estimation hinges on estimating $\|\overline{\mathbf{u}}_{t}\|_{2}^{2}$ efficiently.
Upcoming sections will detail two techniques for communication efficient variance estimation (which primarily involves estimating $\|\overline{\mathbf{u}}_{t}\|_{2}^{2}$): SketchFDA and LinearFDA. To present them uniformly, we introduce the local state $\mathbf{S}_{t}^{(k)}$, a tensor which contains: (1) the scalar value $\|\mathbf{u}^{(k)}_{t}\|_{2}^{2}$ for precisely calculating the first quantity, and (2) a low-dimensional summary of $\mathbf{u}_{t}^{(k)}$, different for each technique, for estimating the second quantity. For each technique we define an estimation function $H(\cdot)$ that calculates the current variance estimate from average state $\overline{\mathbf{S}}_{t}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{S}_{t}^{(k)}$ (obtained via AllReduce).